# MTU: The Multifunction Tree Unit in zkSpeed for Accelerating HyperPlonk

### v1

@@ 5 Methodology
Simulation and Workloads. We evaluate MTU using the representative workloads described in Section 3.1. Software performance is measured on Intel Xeon Gold 5218 CPU [24] with Thermal Design Power (TDP) 125 W, and DDR4 memory. The baseline software implementation is built on top of the HyperPlonk and the arkworks Rust library [45, 3] Software multi-threaded tests are handled by Rayon. To explore scalability, we evaluate the CPU baseline with thread counts ranging from 1 to 32, and compare it against MTU configurations using 2 to 32 PEs. Since MTU uses hybrid binary tree traversal with a DFS Accumulator, we start from a configuration with 2 PEs. We do not evaluate Hybrid traversal on the CPU, as it lacks the tight pipeline structure of hardware PEs, making it infeasible for the precise scheduling required for the DFS Accumulator stage in Hybrid traversal. To reflect realistic system memory conditions, we simulate MTU performance under various bandwidth constraints, ranging from 64 GB/s (comparable to CPU systems with DDR4 [29, 9] and DDR5 [32]) to 1024 GB/s (HBM-class memory [23, 1, 28, 25]). This allows us to evaluate MTU’s performance potential across both commodity and high-end memory environments.
Hardware Implementation. We design a cycle-accurate simulator to model MTU’s hybrid traversal, as discussed in Section 4, together with a scheduler to control the task for the DFS accumulator in MTU. Our simulation uses 255-bit datatypes, consistent with the finite field setting in HyperPlonk. We use Catapult HLS 2023 to generate fully-pipelined Montgomery multipliers supporting arbitrary prime moduli (following prior work [14, 13]). For hash computations in Merkle Tree, we instantiate the SHA3 block from OpenCores [30]. On-chip memory is modeled using SRAM estimates generated by Synopsys 22 nm Memory Compiler. Area and power figures are scaled to 7 nm using technology scaling factors: 3.6×3.6$\times$3.6 × for area, $\times3$ for power, and 1.7×1.7$\times$1.7 × for delay [14, 12, 40, 39]. The MTU is assumed to operate at a clock frequency of 1 GHz.
6 Evaluation
In this section we evaluate each workload with different traversal methods on CPU and MTU. We demonstrate the performance scaling on CPU threads, and the hardware performance scaling with constraint bandwidth.
6.1 Traversal Analyses on CPU
We first evaluate various binary tree traversal schemes on a CPU. Figure 4 presents the runtime performance for four representative workloads: constructing the Build MLE (resulting in an MLE table of size $2^{20}$ entries), evaluating the MLE, computing the Product MLE, and performing Merkle Tree commitment, each over an input table of size $2^{20}$. By varying the number of CPU threads, we compare the traversal and analyze how traversal methods benefit from parallelism.
Among the workloads, the Merkle Tree shows clear performance scaling with increased thread count. In contrast, workloads involving modular arithmetic – such as Build MLE, MLE evaluation, and Product MLE – do not scale linearly with threads. This difference stems from the nature of the underlying computation: the SHA3 hash used in the Merkle Tree is well-optimized on modern CPUs via intrinsic SHA Extensions [22] and benefits significantly from thread-level parallelism, whereas modular arithmetic operations are computationally intensive with more computation dependency.
@@ 6.1 Traversal Analyses on CPU
Figure 5: MTU runtime results across different workloads (each size $2^{20}$), grouped by different traversals. Each group shows hardware runtime under varying hardware bandwidths and PE counts. We start from 2 PEs to perform Hybrid traversal.
@@
The performance difference between BFS and DFS traversal is not pronounced, especially for large, compute-heavy workloads like Merkle Tree commitment, where most of the execution time is dominated by computation. The overhead introduced by traversal patterns becomes relatively minor in the overall execution time. Nevertheless, DFS shows a modest runtime advantage (typically under 200 ms). This is attributed to two factors. First, in BFS, each tree level is processed using Rayon’s parallel iterators. While this simplifies parallelization, it introduces scheduling overhead: threads are spawned at each level of the tree, causing latency from synchronization and load balancing. In contrast, DFS implementation statically partitions the binary tree into independent subtrees, each assigned to a thread before the recursion. This allows threads to proceed without interruption, thus improving efficiency. Second, DFS achieves better cache locality compared to BFS. As discussed before, BFS traversal has fewer local reuse and accesses cross-level caches to store a large amount of data. In contrast, DFS naturally promotes sequential and localized memory access due to its recursive structure, allowing intermediate data to be consumed quickly. This increases the likelihood of cache hits and leads to more efficient execution for large workloads. While these advantages do not drastically change the overall asymptotic runtime, they could contribute to a consistent latency difference in practice.
Overall, the performance difference between BFS and DFS is minor for these workloads, as the computations are already dominated by intensive modular arithmetic or cryptographic hashing. In such compute-bound scenarios and large-size workloads, the traversal strategy contributes only marginally to total runtime. This observation suggests that traversal choices may matter more in memory- or bandwidth-bound contexts, especially in hardware accelerators where compute-latency is greatly improved.
6.2 MTU Parallel Execution and Speedup
Figure 6: Hardware speedup over CPU baseline on workloads of size $2^{20}$ each, grouped by different traversals. Each group shows hardware performance under two bandwidth settings: DDR-level (low) and HBM-level (high). We start from 2 PEs to perform Hybrid traversal.
@@ 6.2 MTU Parallel Execution and Speedup
We now evaluate the performance of the MTU hardware design across different numbers of processing elements (PEs), traversal strategies, and memory bandwidth constraints, ranging from DDR-level (64 GB/s) to HBM-scale (1024 GB/s). As in Section 6.1, we consider four representative ZKP workloads: Build MLE, MLE evaluation, Product MLE, and Merkle Tree commitment, each with a workload size of $2^{20}$. Figure 5 shows the runtime results under these varying configurations. For the BFS traversal, each PE operates independently on nodes of the same tree level. In contrast, in the DFS traversal, each PE processes a disjoint subtree (as illustrated in Figure 1 and 2), and a final PE merges the results. The MTU’s Hybrid traversal, combines BFS and DFS to balance memory reuse and parallelism. We start from two PEs as it needs at least two PEs to build Hybrid traversal in MTU.
Traversal Comparison. We first analyze the difference between traversal methods. Under constrained bandwidth (e.g., 64 GB/s) and low parallelism (2 PEs), both DFS and Hybrid traversals significantly outperform BFS. This is because BFS reads and writes each tree level to off-chip memory, making it heavily bandwidth-bound even with a single PE. In contrast, DFS and Hybrid traversals retain intermediate values on-chip and only access off-chip memory for the initial input (or final output in the case of Build MLE), avoiding bandwidth-induced stalls. An exception is the Product MLE workload, where it takes every tree level (including the intermediate levels) as output results. Therefore, even in DFS and Hybrid modes, performance becomes bandwidth-limited due to the increased off-chip traffic for outputs storing. On the other hand, in a high-bandwidth regime (1024 GB/s) with 2 PEs, traversal choice makes little difference, as all variants become compute-bound by the PE. For workloads except Product MLE, the DFS and Hybrid traversals achieve nearly 3$\times$ speedup over BFS, with sufficient bandwidth and compute resources.
@@
Scalability with PE Parallelism. The difference between traversals can be further evaluated by scaling the number of PEs under both bandwidth constraints. At 64 GB/s, increasing the PE count yields little benefit across all traversal methods due to memory bandwidth saturation. With 1024 GB/s, the performance scales with PE count; however, BFS saturates quickly – plateauing after 16 PEs – because of its bandwidth-heavy nature. Product MLE workload exhibits similar limitations after 16 PEs, even under DFS or Hybrid traversal, since its per-level output demands high sustained bandwidth. For the other workloads, DFS and Hybrid traversals scale more effectively beyond 16 PEs, demonstrating better bandwidth efficiency.
@@
Although DFS and Hybrid deliver comparable performance, it is important to note a key limitation of DFS: it assumes the input is divided into disjoint subtrees. This assumption makes DFS unsuitable when the upstream component delivers inputs in a contiguous, streaming fashion. In contrast, Hybrid traversal supports both parallelism and continuous input indexing, making it more practical in pipeline-integrated systems.
@@
Our evaluation also reveals the significant impact of memory bandwidth on performance scalability. With 32 PEs, the system becomes bandwidth-bound – performance is limited unless sufficient memory bandwidth is provisioned. As shown in Figure 5, increasing the bandwidth from 64 GB/s to 1024 GB/s enables near-linear scaling, as the additional bandwidth alleviates the data movement bottleneck and allows the hardware to maintain high utilization across all PEs.
Speedup over CPU. We compare the performance of MTU against the CPU baseline under matched PE/thread configurations. For the CPU, we select the best-performing traversal method as the baseline. On the hardware side, we evaluate MTU at two bandwidth levels: 64 GB/s, which is comparable to typical DDR4 memory bandwidth on CPUs, and 1024 GB/s, representing HBM-class bandwidth to illustrate the performance ceiling of MTU. Speedup results are shown in Figure 6.
@@
The observed speedup trends mirror the hardware runtime patterns as discussed earlier. Notably, under the 64 GB/s bandwidth constraint, MTU’s speedup diminishes as the number of PEs increases. This is because the hardware falls bandwidth-bound, limiting performance gains despite additional PEs. Meanwhile, CPU performance continues to scale with increasing threads, as the workload remains compute-bound (as shown in Figure 4). Consequently, the relative hardware speedup decreases when CPU runtime improves with additional threads.
@@
Overall, under CPU-level bandwidth conditions (64 GB/s), MTU with Hybrid traversal achieves an average speedup of 1478×1478$\times$1478 × over the CPU baseline. When provisioned with high bandwidth (1024 GB/s), MTU can scale significantly better, reaching up to 9440×9440$\times$9440 × average speedup against CPU due to its efficient high parallelism and sufficient memory bandwidth.
6.3 Area, Power and Pareto Space Analysis
@@
Figure 7: Runtime-area design space of MTU using Hybrid traversal on the Merkle Tree Commit workload. Each point represents a design configuration with a number of PEs and bandwidth. The Pareto frontier illustrates the trade-off between runtime and area across different bandwidth settings.
@@
Prospect of MTU. Table 4 summarizes the area and thermal design power of a representative MTU configuration with 32 PEs. MTU exhibits a compact footprint, and its area can be further reduced under limited bandwidth settings by scaling down the number of PEs. This enables MTU to be embedded within a larger SoC or included as part of a chiplet-based system while sharing a common HBM PHY interface for off-chip memory access. For example, MTU is integrated into the zkSpeed accelerator [13] to support ZKP proof generation. Beyond this case, MTU can be deployed to support potential SumCheck accelerators for MLE-related computation, or repurposed as a polynomial commitment engine within broader ZKP hardware stacks. Its versatility and compactness make MTU a promising building block for modular ZKP accelerator architectures across a range of proof systems and deployment platforms.
Table 4: Area and Thermal Design Power (TDP) breakdown of MTU hardware configured with 32 PEs.
@@
7 Related work
Several prior works have explored hardware acceleration for ZKP protocols across various platforms and protocol families. SZKP [14] accelerates the Groth16 protocol by optimizing MSM using Pippenger’s algorithm and dedicated scheduling strategies. MSM over elliptic curves serves as a PCS alternative to Merkle Tree-based commitments, offering smaller proof sizes but at higher computational cost. NoCap [31] proposes a vector-based processor targeting Spartan+Orion ZKP protocols. These protocols rely on SumCheck and Merkle Tree as core building blocks. While NoCap supports SumCheck, it does not provide specialization for SumCheck subroutines such as Build MLE or MLE evaluation. Merkle Tree computation is mapped to the vector engine directly using level-order BFS traversal. However, NoCap uses 64-bit datatypes while our implementation targets 255-bit. zkSpeed [13] is an ASIC accelerator that targets HyperPlonk. It performs end-to-end acceleration of proof generation, with dedicated modules for MSM, SumCheck and its associated subroutines. UniZK [38] presents a systolic-array-based architecture that accelerates NTT and Merkle Tree. It optimized the Merkle Tree by partitioning a tree into subtrees, and executing each subtree on-chip with BFS. UniZK uses Poseidon hash, whereas both NoCap and MTU use SHA3.
Several prior works have explored traversal optimizations to address application-specific constraints. Optimizations in [26, 5] aim to improve efficiency when the total workload size is unknown and reduce redundant hash computations by speculating on future authentication paths. Other hardware-focused work [35] improves Merkle Tree updates by reducing unnecessary memory traffic through cache-aware designs. For non-cryptographic workloads, for example, [2] proposes switching between BFS and DFS during Random Forest training depending on the active sample size. While these optimizations are effective for their target applications, a unified hardware evaluation of traversal strategies is not provided for the diverse tree-based workloads in modern ZKP protocols.
8 Conclusion
We presents a systematic evaluation of binary tree-based workloads in ZKP, focusing on traversal strategies and hardware design trade-offs on MTU. Our results demonstrate that traversal strategies significantly impact performance, and that the proposed Hybrid Traversal achieves substantial improvements in both runtime and memory efficiency. Using the MTU architecture, we show up to 1478$\times$ speedup over CPU baselines under DDR-level bandwidth. These findings offer practical insights for future ZKP accelerator designs and highlight the potential of tree units like MTU in modular ZKP hardware stacks. Its generality, efficiency, and portability makes MTU a promising building block for modular ZKP hardware platforms.

### v2

@@ 5. Methodology
Simulation and Workloads. We evaluate MTU using the representative workloads described in Section 3.1. Software performance is measured on Intel Xeon Gold 5218 CPU (Intel, 2025) with Thermal Design Power (TDP) 125 W, and DDR4 memory. The baseline software implementation is built on top of the HyperPlonk and the arkworks Rust library (Zhang et al., 2022; arkworks, 2022). Software multi-threaded tests are handled by Rayon. To explore scalability, we evaluate the CPU baseline with thread counts ranging from 1 to 32, and compare it against MTU configurations using 2 to 32 PEs. Since MTU uses hybrid binary tree traversal with a DFS Accumulator, we start from a configuration with 2 PEs. We do not evaluate Hybrid traversal on the CPU, as it lacks the tight pipeline structure of hardware PEs, making it infeasible for the precise scheduling required for the DFS Accumulator stage in Hybrid traversal. To reflect realistic system memory conditions, we simulate MTU performance under various bandwidth constraints, ranging from 64 GB/s (comparable to CPU systems with DDR4 (Nukada, 2021; by Micron Technology Inc., 2023), DDR5 (Schlachter and Drake, 2019)) to 1024 GB/s (HBM-class memory (Inc., 2020; hbm, 2023; Mo et al., 2023b; Kim et al., 2021; Mo and Reagen, 2023)). This allows us to evaluate MTU’s performance potential across both commodity and high-end memories.
@@
Figure 4. MTU runtime across workloads of size $2^{20}$, grouped by traversal types. Each group shows hardware runtime under varying hardware bandwidths (GB/s) and PE counts. Note that DFS is parallelized by partitioning the binary tree into subtrees, resulting in non-contiguous input/output indices.
Hardware Implementation. We design a cycle-accurate simulator to model MTU’s hybrid traversal, as discussed in Section 4, together with a scheduler to control the task for the DFS accumulator in MTU. Our simulation uses 255-bit datatypes, consistent with the finite field setting in HyperPlonk. We use Catapult HLS 2024 to generate fully-pipelined Montgomery multipliers supporting arbitrary prime moduli (following prior work (Daftardar et al., 2024, 2025b)). For hash computations in Merkle Tree, we instantiate the SHA3 block from OpenCores (OpenCores, 2013). On-chip memory is modeled using SRAM generated by Synopsys 22 nm Memory Compiler. Area and power figures are scaled to 7 nm using technology scaling factors: $3.6\times$ for area, $3.3\times$ for power, and $1.7\times$ for delay (Daftardar et al., 2024; Company, 2021; Wu et al., 2016, 2013; Geng et al., 2023). The MTU is assumed to operate at a clock frequency of 1 GHz.
Figure 5. CPU performance across different workloads using BFS and DFS. Each workload is evaluated at size $2^{20}$.
6. Evaluation
Here, we evaluate each workload with different traversal methods on CPU and MTU. We show performance scaling on CPU threads and hardware performance scaling with constrained bandwidth.
6.1. Traversal Analyses on CPU
We first evaluate various binary tree traversal schemes on a CPU. Figure 5 presents the runtime performance for four representative workloads: constructing the Build MLE (resulting in an MLE table of size $2^{20}$ entries), evaluating the MLE, computing the Product MLE, and performing Merkle Tree commitment, each over an input table of size $2^{20}$. Under varying number of CPU threads, we compare traversal schemes and analyze how they benefit from parallelism.
Among the workloads, the Merkle Tree shows clear performance scaling with increased thread count. In contrast, workloads involving modular arithmetic – such as Build MLE, MLE evaluation, and Product MLE – do not scale linearly with threads. This difference stems from the nature of the underlying computation: the SHA3 hash used in the Merkle Tree is well-optimized on modern CPUs via intrinsic SHA Extensions (Gully et al., 2013) and benefits significantly from thread-level parallelism, whereas modular arithmetic operations are computationally intensive with more computation dependency.
@@ 6.1. Traversal Analyses on CPU
The performance difference between BFS and DFS traversal is not pronounced, especially for large, compute-heavy workloads like Merkle Tree commitment, where most of the execution time is dominated by computation. The overhead introduced by traversal patterns becomes relatively minor in the overall execution time. Nevertheless, DFS shows a modest runtime advantage (typically under 200 ms). This is attributed to two factors. First, in BFS, each tree level is processed using Rayon’s parallel iterators. While this simplifies parallelization, it introduces scheduling overhead: threads are spawned at each level of the tree, causing latency from synchronization and load balancing. In contrast, DFS implementation statically partitions the binary tree into independent subtrees, each assigned to a thread before the recursion. This allows threads to proceed without interruption, thus improving efficiency. Second, DFS achieves better cache locality compared to BFS. As discussed before, BFS traversal has less local reuse and accesses cross-level caches to store a large amount of data. In contrast, DFS naturally promotes sequential and localized memory access due to its recursive structure, allowing intermediate data to be consumed quickly. This increases the likelihood of cache hits and leads to more efficient execution for large workloads. While these advantages do not drastically change the overall asymptotic runtime, they could contribute to a consistent latency difference in practice.
Overall, in such compute-bound scenarios, traversal strategy has limited impact. It matters more in the memory- or bandwidth-bound cases, particularly with hardware accelerators.
6.2. MTU Parallel Execution and Speedup
@@
We now evaluate MTU performance across different numbers of processing elements (PEs), traversal strategies, and memory bandwidth constraints, ranging from DDR-level (64 GB/s) to HBM-scale (1024 GB/s). As in Section 6.1, we consider four representative ZKP workloads: Build MLE, MLE evaluation, Product MLE, and Merkle Tree commitment, each with a workload size of $2^{20}$. Figure 4 shows the runtime results under these varying configurations. For the BFS traversal, each PE operates independently on nodes of the same tree level. In contrast, in the DFS traversal, each PE processes a disjoint subtree (see Figure 1 and 2), and a final PE merges the results. The MTU’s Hybrid traversal combines BFS and DFS to balance memory reuse and parallelism. We start from two PEs as the MTU needs at least two PEs to support Hybrid traversal.
[…]
Figure 6. Hardware speedup over CPU baseline on workloads (each size $2^{20}$), grouped by different traversals. Each group shows DDR (low) and HBM-level (high) bandwidth (GB/s). We start from 2 PEs to perform Hybrid traversal.
@@
Scalability with PE Parallelism. The difference between traversals can be further evaluated by scaling the number of PEs under both bandwidth constraints. At 64 GB/s, increasing the PE count yields little benefit across all traversal methods due to memory bandwidth saturation. With 1024 GB/s, the performance scales with PE count; however, BFS saturates quickly (plateauing after 16 PEs) because of its bandwidth-heavy nature. Product MLE exhibits similar limitations after 16 PEs, even under DFS or Hybrid traversal, since its per-level output demands high sustained bandwidth. For the other workloads, DFS and Hybrid traversals scale more effectively beyond 16 PEs, exhibiting better bandwidth efficiency.
@@
Although DFS and Hybrid deliver comparable performance, a key limitation of DFS is its assumption that its input is divided into disjoint subtrees. DFS may be unsuitable when upstream modules deliver inputs in a contiguous, streaming fashion. In contrast, Hybrid traversal supports both parallelism and continuous input indexing, making it more practical in pipeline-integrated systems.
@@
Our evaluation also reveals the significant impact of memory bandwidth on performance scalability. With 32 PEs, the system becomes memory-bound – performance is limited unless sufficient bandwidth is provisioned. As shown in Figure 4, increasing the bandwidth from 64 GB/s to 1024 GB/s enables near-linear scaling, as the additional bandwidth alleviates the data movement bottleneck and allows the hardware to maintain high utilization across all PEs.
@@
Speedup over CPU. We compare MTU performance against CPU baselines under matched PE/thread configurations. For the CPU, we select the best-performing traversal method as the baseline. For MTU evaluations, we use two bandwidth levels: 64 GB/s, which is comparable to typical DDR4 memory bandwidth on CPUs, and 1024 GB/s, representing HBM-class bandwidth to illustrate the performance ceiling of MTU. Speedup results are shown in Figure 6. The observed speedup trends mirror the hardware runtime patterns as discussed earlier. Notably, under the 64 GB/s bandwidth constraint, MTU’s speedup diminishes as the number of PEs increases because the hardware is memory-bound. Meanwhile, CPU performance continues to scale with increasing threads, as the workload remains compute-bound (see Figure 5), so the relative hardware speedup decreases as the CPU improves with additional threads.
@@
Overall, under CPU-class bandwidths (64 GB/s), MTU with Hybrid traversal achieves an average speedup of $1478\times$ over the CPU baseline. Under higher bandwidths (1024 GB/s), MTU scales much better, reaching up to $9440\times$ average speedup because it can exploit high degrees of parallelism and sufficient memory bandwidth.
@@
Figure 7. Runtime-area design space of MTU using Hybrid traversal on the Merkle Tree Commit workload. Each point represents a design configuration with a number of PEs and bandwidth. The Pareto frontier illustrates the trade-off between runtime and area across different bandwidth settings.
6.3. Area, Power and Pareto Space Analysis
@@
Prospect of MTU. Table 4 summarizes the area and thermal design power of a representative MTU configuration with 32 PEs. MTU exhibits a compact footprint, and its area can be further reduced under limited bandwidth settings by scaling down the number of PEs. This enables MTU to be embedded within a larger SoC or included as part of a chiplet-based system while sharing a common HBM PHY interface for off-chip memory access. For example, MTU is integrated into the zkSpeed accelerator (Daftardar et al., 2025b) to support ZKP generation. Beyond this case, MTU can be deployed to support potential SumCheck accelerators for MLE-related computation, or repurposed as a polynomial commitment engine within broader ZKP hardware stacks. Its versatility and compactness make MTU a promising building block for modular ZKP accelerator architectures across a range of proof systems and deployment platforms.
7. Related work
Several prior works have explored hardware acceleration for ZKP protocols across various platforms and protocol families. SZKP (Daftardar et al., 2024) accelerates the Groth16 protocol by optimizing MSM using Pippenger’s algorithm and dedicated scheduling strategies. MSM over elliptic curves serves as a PCS alternative to Merkle Tree-based commitments, offering smaller proof sizes but at higher computational cost. NoCap (Samardzic et al., 2024) proposes a vector-based processor targeting Spartan+Orion ZKP protocols. These protocols rely on SumCheck and Merkle Tree as core building blocks. While NoCap supports SumCheck, it does not provide specialization for SumCheck subroutines such as Build MLE or MLE evaluation. Merkle Tree computation is mapped to the vector engine directly using level-order BFS traversal. zkSpeed (Daftardar et al., 2025b) is an ASIC that accelerates HyperPlonk proof generation end-to-end, with dedicated modules for MSM, SumCheck and its associated subroutines. UniZK (Wang and Gao, 2025) presents a systolic-array-based architecture that accelerates NTT and Merkle Tree. It optimized the Merkle Tree by partitioning a tree into subtrees, and executing each subtree on-chip with BFS.
Table 4. Area/power breakdown of MTU hardware (32 PEs). Misc includes the control and scheduler cost.
@@ 7. Related work
Several prior works have explored traversal optimizations to address application-specific constraints. Optimizations in (Knecht and Nicola, 2013; Berman et al., 2007) aim to improve efficiency when the total workload size is unknown and reduce redundant hash computations by speculating on future authentication paths. Other hardware-focused work (Shadab et al., 2022) improves Merkle Tree updates by reducing unnecessary memory traffic through cache-aware designs. For non-cryptographic workloads, for example, (Anghel et al., 2019) proposes switching between BFS and DFS during Random Forest training depending on the active sample size. While these optimizations are effective for their target applications, a unified hardware evaluation of traversal strategies is not provided for the diverse tree-based workloads in modern ZKP protocols.
8. Conclusion
[…]
Acknowledgements.
This research was developed with funding in part from the NSF CAREER award #2340137, the NSF CIRC Grand award #2450539, and from DARPA under the Hybrid Electromagnetic side-channel and Interactive-proof Methods to Detect and Amend LogicaL Rifts (HEIMDALLR) program, grant number HR0011-25-C-0300.